# Selective Transformer for Hyperspectral Image Classification

### v2

@@ Abstract
Transformer has achieved satisfactory results in the field of hyperspectral image (HSI) classification. However, existing Transformer models face two key challenges when dealing with HSI scenes characterized by diverse land cover types and rich spectral information: (1) fixed receptive field representation overlooks effective contextual information; (2) redundant self-attention feature representation. To address these limitations, we propose a novel Selective Transformer (SFormer) for HSI classification. The SFormer is designed to dynamically select receptive fields for capturing both spatial and spectral contextual information, while mitigating the impact of redundant data by prioritizing the most relevant features. This enables a highly accurate classification of the land covers of the HSI. Specifically, a Kernel Selective Transformer Block (KSTB) is first utilized to dynamically select an appropriate receptive field range to effectively extract spatial-spectral features. Furthermore, to capture the most crucial tokens, a Token Selective Transformer Block (TSTB) is introduced, which selects the most relevant tokens based on the ranking of attention scores for each query. Extensive experiments on four benchmark HSI datasets demonstrate that the proposed SFormer outperforms the state-of-the-art HSI classification models. The codes will be released.
Index Terms:
Hyperspectral image (HSI) classification, transformer, selective spatial attention, fine-grained spectral attention.
I Introduction
@@
With the continuous advancements in hyperspectral imaging technology, hyperspectral images (HSIs) now offer increasingly rich spatial-spectral information, enabling precise Earth observation[1, 2, 3, 4]. Leveraging their distinctive spectral characteristics, HSIs are capable of performing a wide range of tasks, including classification[5, 6, 7], target detection[8, 9], change detection[10, 11], and image quality enhancement[12, 13]. Notably, HSI classification has become a fundamental task in the remote sensing community due to its extensive applications in areas such as disaster monitoring[14], precision agriculture[15], and urban planning[16]. The primary goal of HSI classification is to assign specific class labels to each pixel within the HSI.
Deep neural networks have become dominant in the field of HSI classification, including stacked auto-encoders (SAEs)[17], deep belief networks (DBNs)[18], recurrent neural networks (RNNs)[19], graph convolutional networks (GCNs)[20], convolutional neural networks (CNNs)[21], vision transformers (ViTs)[22] and recently emerging state space models (SSMs)[23]. Deep learning methods offer the advantage of automatically learning discriminative deep features from input images without relying on hand-crafted features.
Owing to their powerful feature representation capabilities, CNNs[24, 25, 26] have attracted significant attention. A 1D-CNN[27] was proposed to achieve efficient local spectral feature extraction through multiple convolutional layers. Further, Zhao et al. [28] introduced 2D-CNN that extracts spatial-spectral features and uses a balanced local discriminant embedding. Later, 3D-CNNs [29, 30] were developed to directly extract joint spatial-spectral features from the HSI cube. Then, some multiscale convolutional networks[31, 32] were proposed to extract deep multiscale features. However, these CNN-based networks are constrained by their inherent local perception, which limits the amount of information they can capture, and thus hinders accuracy.
Subsequently, HSI classification approaches with attention mechanism[33, 34, 35, 36] emerged and achieved significant classification results, particularly self-attention mechanisms[37, 38, 39], due to their ability to improve long-range perception. After that, transformer-based networks[40, 41, 42, 22, 43, 44, 45, 46] have been widely adopted for their use of Multi-Head Self-Attention (MHSA) to achieve superior context modeling. These approaches have notably outperformed traditional CNNs, which are limited by their receptive fields. Hong et al. [22] proposed SpectralFormer, a network that enhances HSI classification by capturing spectral sequence information between adjacent HSI bands using transformers and cross-layer skip connections. Further, Sun et al. [43] proposed a spectral–spatial feature tokenization transformer (SSFTT), where shallow features are extracted by convolutions, semantic features are processed by a Gaussian weighted tokenization module, and a single-layer transformer is used for feature representation. Then, Zou et al. [47] introduced LESSFormer, which employs an HSI2Token module to extract adaptive spatial-spectral tokens and uses a local transformer to enhance the local spectral representation. Inspired by the effectiveness of spectral partitioning, Mei et al. [48] introduced a grouped pixel embedding strategy and hierarchical structure to extract discriminative multiscale spatial-spectral features. Roy et al. [49] incorporated morphological characteristics into their proposed MorphFormer by combining spectral and spatial morphological convolution operations with attention mechanisms to effectively integrate structural and morphological information. After that, Zhao et al. [45] introduced a lightweight variant of the ViT network, which combines group-separable convolution and Multi-Head Self-Attention modules within feature extraction blocks to efficiently capture local and global spatial features with reduced parameter complexity.
@@
However, most transformer structures typically utilize a fixed receptive field for feature extraction, overlooking the fact that different types of land cover require varying contextual information[50, 51]. As shown in Fig. 1 (a), HSI with a limited receptive field may exhibit misclassification issues. Different types of land cover often require varying ranges of contextual information to allow an accurate classification of pixels within the HSI. Moreover, most transformer-based HSI classification models usually use the vanilla Multi-Head Self-Attention to build global features[48]. Fig. 1 (b) illustrates tokens of the land cover category related to the query token, as well as some tokens not related to the type of land cover target. In addition, existing transformer-based HSI classification models predominantly focus on spatial attention or spectral attention[22], neglecting the integrated modeling of spatial and spectral contextual information within HSI data.
@@
To address these issues, this paper proposes a novel Selective Transformer (SFormer) for HSI classification, which selects the most valuable and relevant spatial-spectral information within a reasonable receptive field. SFormer is designed with the understanding that objects of different scales require varying contextual information. To accommodate this, the model incorporates a Kernel Selection Transformer Block (KSTB), which dynamically selects kernels of appropriate sizes, optimizing the receptive fields for effective spatial-spectral feature extraction. Furthermore, rather than utilizing all tokens for dense attention calculation, SFormer introduces a Token Selective Transformer Block (TSTB). This block selects the most relevant tokens with the highest attention values with a group strategy that can directly utilize the 3D HSI characteristic. These selective mechanisms allow SFormer to capture the most significant and appropriate spatial-spectral information within an optimal receptive field. The main contributions of this study are summarized as follows:
Figure 1: (a) An illustration of limited receptive fields in HSI classification shows how varying land-cover types require different scales of context, potentially leading to misclassification, such as confusing grass with bricks. (b) An example of relevant and irrelevant tokens in relation to a query token demonstrates that while relevant tokens provide valuable information for accurately classifying an asphalt land-cover type, irrelevant tokens may introduce noise, potentially impairing classification accuracy.
  • 1)
@@
    This paper proposes a novel Selective Transformer (SFormer) for HSI classification. The SFormer effectively selects receptive fields to extract both spatial and spectral contextual information, while reducing redundant information interference by focusing on the most relevant features. This approach enables the precise classification of HSI ground objects.
  • 2)
@@
    To adaptively select the appropriate contextual information, we design a Kernel Selection Transformer Block (KSTB). This mechanism ensures that feature extraction occurs within an optimal receptive field from both spatial and spectral perspectives, thereby enhancing the model’s capacity to effectively learn and distinguish diverse land cover features.
  • 3)
@@
    We further develop a Token Selective Transformer Block (TSTB) that strategically selects the most relevant tokens to extract valuable information. This mechanism effectively integrates essential information while reducing the impact of redundant and irrelevant information, thereby improving the representation of spatial-spectral features in visual tasks.
  • 4)
@@
    Extensive experiments are conducted on four benchmark HSI datasets: Pavia University, Houston, Indian Pines, and WHU-HongHu. The results indicate that the proposed SFormer, which utilizes appropriate and valuable contextual information, performs better than other state-of-the-art methods.
@@
The remainder of this paper is organized as follows. Section II provides a comprehensive description of the proposed SFormer model. Section III presents and thoroughly analyzes the experimental results. Finally, Section IV concludes the paper.
II Proposed Methods
In this section, we present the implementation details of our SFormer. SFormer consists of convolutional layers, selective transformer groups (STG), and a classification head. We begin with an overview of SFormer, followed by a detailed explanation of its components: the Kernel Selective Transformer Block (KTAB), and the Token Selective Transformer Block (TSTB).
@@ II Proposed Methods
Figure 2: An illustration of the proposed SFormer. The Selective Transformer Group (STG) is composed of a Kernel Selective Transformer Block (KSTB) and three consecutive Token Selective Transformer Blocks (TSTBs). Followed by two STGs, we adopt the fully connected layer as the classification head to obtain the classification results.
II-A Overview of SFormer
@@
Figure 3: The proposed Kernel Selective Attention (KSA) conceptual illustration. Here, DwConv represents Depth-wise convolution, Avg and Max represent channel-wise average pooling and max pooling, respectively, and GAP means spatial global average pooling.
@@
As shown in Fig. 2, the HSI is first subjected to PCA for dimensionality reduction, followed by the extraction of overlapping patch cubes, $\{x_{i}\in\mathbb{R}^{c\times w\times w}\}_{i=1}^{H\times W}$, where each patch is classified based on the central pixel. Here, $H$ and $W$ denote the height and width, $w$ represents the patch size, and $c$ is the reduced spectral dimension. Taking the pixel-level patch cube $x\in\mathbb{R}^{c\times w\times w}$ as input, we first employ a 3×3 convolutional layer to capture shallow features, denoted as $\mathbf{F}_{1}$. Next, we perform deep feature extraction on $\mathbf{F}_{1}$ using two consecutive Selective Transformer Groups (STGs). This process can be formulated as: $\mathbf{F}_{2}=\text{STG}_{2}(\text{STG}_{1}(\mathbf{F}_{1}))$. Finally, a layer normalization followed by a fully connected layer in the classification head is applied to obtain the final classification results.
II-B Kernel Selective Transformer Block
@@
As analyzed in the Introduction, different land cover classes in HSI require distinct contextual information. Existing methods are often limited by the fixed size of the receptive field of the network or cannot dynamically adapt the receptive field to suit specific types of land cover[48]. To address this issue, we designed the Kernel Selective Transformer Block (KSTB). First, it employs dilated depthwise convolutions to construct a larger receptive field for multiscale modeling. Second, it incorporates a spatial-spectral selection mechanism to adaptively determine the appropriate receptive field size required for different types of land cover. Specifically, KSTB primarily consists of a Kernel Selective Attention (KSA) module and a Feed-Forward Network (FFN).
@@
The structure of KSA is shown in Fig. 3, two decoupled depth-wise convolutions (DwConv) are employed for input $\mathbf{X}\in\mathbb{R}^{c\times w\times w}$. The features generated with decomposed kernels are then processed by 1×1 convolution layers $\mathcal{F}_{1\times 1}(\cdot)$ to integrate channel information and unify dimensions, obtaining contextual features with different receptive fields:
@@
The KSA module generates its final output by performing an element-wise multiplication between the input features $\mathbf{X}$ and $\mathbf{S}$:
@@
II-C Token Selective Transformer Block
Previous research has typically utilized the self-attention mechanism within Transformers to capture long-range dependencies, facilitating the modeling of global spatial or spectral relationships[22]. However, HSI often contains a wide variety of complex land cover types. The conventional Transformer architecture tends to compute attention across all query-key pairs, which inevitably leads to interactions between redundant and irrelevant information, ultimately diminishing interpretative accuracy.
Figure 4: The illustration of the proposed Multi-Head Selective Attention (MSA). The input feature $P$ is first grouped and then fed through a 3D Point-wise convolution, followed by a 3D Depth-wise convolution to generate the corresponding $\mathbf{Q}\mathbf{K}\mathbf{V}$. Subsequently, a Token selection mechanism is utilized for further self-attention operation.
To overcome this limitation, we propose a novel multi-head selective attention (MSA) mechanism to replace the traditional multi-head self-attention (MHSA) mechanism. Furthermore, our MSA retains the data characteristics of the HSI three-dimensional cube through a grouped operation and employs 3D convolutions to extract tokens with spatial and spectral properties. During the computation of self-attention, the MSA selectively focuses on the most relevant tokens, effectively disregarding the least relevant information. This approach improves interpretive accuracy while reducing redundant computations.
As illustrated in Fig. 4, we first begin by partitioning the input feature $\mathbf{P}\in\mathbb{R}^{C\times H\times W}$ into $g$ groups, each producing a feature map of size $c\times H\times W$. Then we feed them through a 3D point-wise convolution with a kernel size of $1×1×$ 1 × 1, followed by a 3D depth-wise convolution with a kernel size of $1×3×3$. This process generates the corresponding query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$. To facilitate the calculation of self-attention, we reshape and transpose the tensors $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, adjusting their dimensions by flattening the operation, so each has a shape of $(H\times W\times g)\times c$. These processes can be extended to the multi-head scenario and expressed as follows:

### v3

@@ Abstract
Transformer has achieved satisfactory results in the field of hyperspectral image (HSI) classification. However, existing Transformer models face two key challenges when dealing with HSI scenes characterized by diverse land cover types and rich spectral information: (1) A fixed receptive field overlooks the effective contextual scales required by various HSI objects; (2) invalid self-attention features in context fusion affect model performance. To address these limitations, we propose a novel Dual Selective Fusion Transformer Network (DSFormer) for HSI classification. DSFormer achieves joint spatial and spectral contextual modeling by flexibly selecting and fusing features across different receptive fields, effectively reducing unnecessary information interference by focusing on the most relevant spatial-spectral tokens. Specifically, we design a Kernel Selective Fusion Transformer Block (KSFTB) to learn an optimal receptive field by adaptively fusing spatial and spectral features across different scales, enhancing the model’s ability to accurately identify diverse HSI objects. Additionally, we introduce a Token Selective Fusion Transformer Block (TSFTB), which strategically selects and combines essential tokens during the spatial-spectral self-attention fusion process to capture the most crucial contexts. Extensive experiments conducted on four benchmark HSI datasets demonstrate that the proposed DSFormer significantly improves land cover classification accuracy, outperforming existing state-of-the-art methods. Specifically, DSFormer achieves overall accuracies of 96.59%, 97.66%, 95.17%, and 94.59% in the Pavia University, Houston, Indian Pines, and Whu-HongHu datasets, respectively, reflecting improvements of 3.19%, 1.14%, 0.91%, and 2.80% over the previous model. The code will be available online at https://github.com/YichuXu/DSFormer.
keywords:
[…]
1 Introduction
@@
With the continuous advancements in hyperspectral imaging technology, hyperspectral images (HSIs) now offer increasingly rich spatial-spectral information, enabling precise Earth Observation (Yang et al., 2024; Zhang et al., 2022; Paoletti et al., 2019; Xu et al., 2023; Wang et al., 2024). Leveraging their distinctive spectral characteristics, HSIs are capable of performing a wide range of tasks, including classification (Shi et al., 2023; Li et al., 2019a; Gao et al., 2024; Roy et al., 2023a), target detection (Zhu et al., 2023; Xu et al., 2022; Xie et al., 2021), change detection (Chang and Ghamisi, 2023; Hu et al., 2023; Chen et al., 2024), and image quality enhancement (Xiao et al., 2024a; Wang et al., 2022b; Chen et al., 2023a). Notably, HSI classification has become a fundamental task in the remote sensing community due to its extensive applications in areas such as disaster monitoring (Hong et al., 2021), precision agriculture (Zhang et al., 2016), and urban planning (van der Linden and Hostert, 2009). The primary goal of HSI classification is to assign specific class labels to each pixel within the HSI.
HSI classification increasingly relies on deep learning techniques, which offer significant advantages over traditional machine learning methods due to their powerful feature extraction capabilities (Zhao and Du, 2016b; Mou et al., 2018). In particularly, convolutional neural networks (CNNs) enable the extraction of features in a hierarchical manner (Roy et al., 2020; Xu et al., 2018a). However, their local convolutional structure limits their ability to capture long-range dependencies. To address this, recent studies (Wang et al., 2018) introduced self-attention mechanisms to capture non-local contextual information, though at a high computational cost due to extensive similarity calculations.
Recently, the success of self-attention in natural language processing (Vaswani et al., 2017) has inspired its application in computer vision. Dosovitskiy et al. (Dosovitskiy et al., 2021) proposed the Vision Transformer (ViT), which segments images into tokens and applies self-attention to capture long-range dependencies, achieving notable improvements in image classification. Transformer-based approaches (He et al., 2024; Zhong et al., 2022; Hong et al., 2022; Zhao et al., 2024) have become a mainstream trend in HSI classification research, with numerous studies introducing novel methods to advance HSI classification technology.
@@
However, most transformer structures typically utilize a fixed receptive field for feature extraction, overlooking the fact that different types of land cover require varying contextual information (Li et al., 2019b, 2023). As shown in Fig. 1 (a), HSI with a fixed receptive field may exhibit misclassification issues. Therefore, we argue that different objects in HSIs should be recognized with correspondingly required receptive fields. After determining the scope of visual perception, the next step is to perform effective context capturing. To achieve this, most transformer-based HSI classification models usually employ vanilla Multi-Head Self-Attention (MHSA) to build global features (Mei et al., 2022) by enabling interactions between one token and all the other tokens. However, since not all tokens are informative, the original MHSA inevitably introduces unexpected noise during context feature extraction. As shown in Fig. 1 (b), the green boxes represent unnecessary tokens relative to the red query marker for asphalt, such as meadows and bitumen. Additionally, in capturing contextual information, existing transformer-based HSI classification models predominantly focus on either spatial or spectral information alone (Hong et al., 2022), neglecting the integrated modeling of spatial-spectral contexts within HSI data.
@@
To address these issues, this paper proposes a novel Dual Selective Fusion Transformer Network (DSFormer) for HSI classification, which selects and fuses the most valuable and relevant spatial-spectral information within a reasonable receptive field. To accommodate this, DSFormer incorporates a Kernel Selective Fusion Transformer Block (KSFTB), which dynamically selects and integrates multiple convolutional kernels of appropriate sizes, optimizing the receptive fields to obtain effective spatial-spectral features. Furthermore, rather than utilizing all tokens for dense attention calculation, DSFormer introduces a Token Selective Fusion Transformer Block (TSFTB). This block selects the most relevant tokens of the highest attention values for fusion through a spectral grouping strategy that can directly utilize the 3D HSI characteristic. These selective mechanisms allow DSFormer to produce the most significant and appropriate spatial-spectral information within an optimal receptive field. The main contributions of this study are summarized as follows:
  1. (1)
@@
    This paper proposes a novel Dual Selective Fusion Transformer Network (DSFormer) for HSI classification. It adaptively selects and fuses features from diverse receptive fields to achieve joint spatial-spectral context modeling, while reducing unnecessary information interference by focusing on the most relevant spatial-spectral tokens. This approach enables the precise classification of HSI ground objects.
  2. (2)
@@
    To adaptively select and integrate the appropriate contextual information, we design a Kernel Selective Fusion Transformer Block (KSFTB). This mechanism ensures that features are acquired within an optimal receptive field from both spatial and spectral perspectives, thereby enhancing the model’s capacity to effectively perceive and distinguish various HSI objects.
  3. (3)
@@
    We further develop a Token Selective Fusion Transformer Block (TSFTB) that strategically selects the most relevant tokens in the context fusion process, effectively integrating essential information while reducing the impact of irrelevant objects, thereby improving the distinguishability of the generated spatial-spectral features for HSI visual interpretation.
  4. (4)
@@
    Extensive experiments are conducted on four benchmark HSI datasets: Pavia University, Houston, Indian Pines, and Whu-HongHu. The results indicate that the proposed DSFormer, which utilizes appropriate and valuable contextual information, performs better than other state-of-the-art methods.
@@
The remaining section of this paper is organized as follows. Section 2 reviews the related work of HSI classification. In Section 3 provides a comprehensive description of the proposed DSFormer model. Section 4 presents and thoroughly analyzes the experimental results. Finally, Section 5 concludes the paper.
Figure 1: The challenges existing in current HSI classification methods. (a) Fixed and limited receptive fields in HSI classification can potentially lead to misclassifications, such as confusing asphalt to bricks. (b) Irrelevant tokens may introduce noise which may potentially impair classification accuracy.
2 Related Works
2.1 CNN-based Methods
With the rapid progress in deep learning, CNNs have shown remarkable potential and significant practical value across a wide range of computer vision tasks, particularly in the field of HSI classification. Owing to their powerful feature representation capabilities, CNNs (Zhao and Du, 2016a; Zhu et al., 2017; Chen et al., 2016; Xu et al., 2020) have attracted significant attention. A 1D-CNN (Hu et al., 2015) was proposed to achieve efficient local spectral feature extraction through multiple convolutional layers. Further, 2D-CNN was introduced in (Zhao and Du, 2016b) that extracts spatial-spectral features and uses a balanced local discriminant embedding. Later, 3D-CNNs (Ben Hamida et al., 2018; Li et al., 2017) were developed to directly extract joint spatial-spectral features from the HSI cube. Then, some multiscale convolutional networks (Xu et al., 2018b; Gong et al., 2019; Wang et al., 2021; Yang et al., 2021) were proposed to extract deep multiscale features.
2.2 Transformer-based Methods
However, the above CNN-based networks are constrained by their inherent local perception, which limits the amount of information they can capture, and thus hinders accuracy. Subsequently, HSI classification approaches with attention mechanism (Sun et al., 2020; Hang et al., 2021; Zhu et al., 2021; Wang et al., 2022a) emerged and achieved significant classification results, particularly self-attention mechanisms (Vaswani et al., 2017; Xu et al., 2021, 2024), due to their ability to improve long-range perception.
@@ 2.2 Transformer-based Methods
Figure 2: An illustration of the proposed DSFormer. The Dual Selective Fusion Transformer Group (DSFTG) is composed of a Kernel Selective Fusion Transformer Block (KSFTB) and three consecutive Token Selective Fusion Transformer Blocks (TSFTBs). Followed by two DSFTGs, we adopt the fully connected layer as the classification head to obtain the classification results.
Transformer-based networks (Chen et al., 2023b; Xiao et al., 2024b; Sun et al., 2022; Cheng et al., 2024; Sun et al., 2024) have been widely adopted for their use of MHSA to achieve superior context modeling. These approaches have notably outperformed traditional CNNs, which are limited by their receptive fields. SpectralFormer was proposed in (Hong et al., 2022), a network that enhances HSI classification by capturing spectral sequence information between adjacent HSI bands using transformers and cross-layer skip connections. Further, Sun et al. (Sun et al., 2022) proposed a spectral–spatial feature tokenization transformer (SSFTT), where shallow features are extracted by convolutions, semantic features are processed by a Gaussian weighted tokenization module, and a single-layer transformer is used for feature representation. Then, LESSFormer was introduced (Zou et al., 2022), which employs an HSI2Token module to extract adaptive spatial-spectral tokens and uses a local transformer to enhance the local spectral representation. Inspired by the effectiveness of spectral partitioning, Mei introduced a grouped pixel embedding strategy and hierarchical structure to extract discriminative multiscale spatial-spectral features (Mei et al., 2022). Yang et al. (Yang et al., 2023) introduced a transformer-based center-to-surround interactive learning (CSIL) framework for extracting multi-scale features for HSI classification tasks. Roy et al. (Roy et al., 2023b) incorporated morphological characteristics into their proposed MorphFormer by combining spectral and spatial morphological convolution operations with attention mechanisms to effectively integrate structural and morphological information. After that, Zhao et al. (Zhao et al., 2024) introduced a lightweight variant of the ViT network, which combines group-separable convolution and MHSA modules within feature extraction blocks to efficiently capture local and global spatial features with reduced parameter complexity.
Compared with the aforementioned Transformer-based methods, we contend that the proposed DSFormer is distinct. First, KSFTB is specifically designed to adaptively choose and fuse relevant context scopes. This mechanism ensures that feature extraction takes place within an optimal receptive field, encompassing both spatial and spectral dimensions. We further develop TSFTB, which aims to combine the most relevant tokens to capture valuable information. This mechanism effectively integrates essential information while mitigating the influence of redundant and irrelevant ingredients, thereby enhancing the representation capability of spatial-spectral features for HSI classification.
3 Methodology
In this section, we present the implementation details of our DSFormer. DSFormer consists of convolutional layers, dual selective fusion transformer groups (DSFTGs), and a classification head. We begin with an overview of DSFormer, followed by a detailed explanation of two components KSFTB, and TSFTB.
@@ 3 Methodology
Figure 3: The proposed Kernel Selective Fusion Attention (KSFA). Here, DwConv represents Depth-wise convolution, Avg and Max represent channel-wise average pooling and max pooling, respectively, and GAP means spatial global average pooling.
3.1 Overview of DSFormer
@@
As shown in Fig. 2, the HSI is first subjected to PCA for dimensionality reduction, followed by the extraction of overlapping patch cubes, $\{x_{i}\in\mathbb{R}^{c\times w\times w}\}_{i=1}^{H\times W}$, where each patch is classified based on the central pixel. Here, $H$ and $W$ denote the height and width, $w$ represents the patch size, and $c$ is the reduced spectral dimension. Taking the pixel-level patch cube $x\in\mathbb{R}^{c\times w\times w}$ as input, we first employ a 3×3 convolutional layer to capture shallow features, denoted as $\mathbf{F}_{1}$. Next, we perform deep feature generation on $\mathbf{F}_{1}$ using two consecutive DSFTGs. This process can be formulated as: $\mathbf{F}_{2}=\text{DSFTG}_{2}(\text{DSFTG}_{1}(\mathbf{F}_{1}))$. Finally, a layer normalization followed by a fully connected layer in the classification head is applied to obtain the final classification results.
3.2 Kernel Selective Fusion Transformer Block
@@
As analyzed in the Introduction, different land cover classes in HSI require distinct contextual scales. Existing methods are often limited by the fixed and limited size of the receptive field of the network or cannot dynamically adapt the receptive field to suit specific types of land cover Sun et al. (2022); Roy et al. (2023b). To address this issue, we designed the KSFTB which learns an optimal receptive field through selecting and integrating multiscale features. KSFTB initially employs dilated depthwise convolutions to construct a large receptive field to obtain sufficient range of feature perception. Then, it incorporates a spatial-spectral selection mechanism to adaptively fusion the appropriate receptive field size required for different types of land cover. Specifically, KSFTB primarily consists of a Kernel Selective Fusion Attention (KSFA) module and a Feed-Forward Network (FFN).
@@
The structure of KSFA is shown in Fig. 3, two decoupled depth-wise convolutions (DwConv) are employed for input $\mathbf{X}\in\mathbb{R}^{c\times w\times w}$. The features generated with decomposed kernels are then processed by 1×1 convolution layers $\mathcal{F}_{1\times 1}(\cdot)$ to integrate channel information and unify dimensions, obtaining contextual features with different receptive fields:
@@
The KSFA module generates its final output by performing an element-wise multiplication between the input features $\mathbf{X}$ and $\mathbf{S}$:
@@
3.3 Token Selective Fusion Transformer Block
Figure 4: The illustration of the proposed Token Selective Fusion Attention (TSFA). The input feature $P$ is first grouped and then fed through a 3D point-wise convolution, followed by a 3D depth-wise convolution to generate the corresponding $\mathbf{Q}\mathbf{K}\mathbf{V}$. Subsequently, a token selection mechanism is utilized for further self-attention operation.
Previous research has typically utilized the self-attention mechanism within Transformers to capture long-range dependencies, facilitating the modeling of global spatial or spectral relationships Hong et al. (2022); Zhang et al. (2024). However, HSI often contains a wide variety of complex land cover types. The conventional Transformer architecture tends to compute attention across all query-key pairs, which inevitably leads to interactions between redundant and irrelevant information, ultimately diminishing interpretative accuracy.
To overcome this limitation, we propose a novel multi-head Token Selective Fusion Attention (TSFA) mechanism to replace the traditional MHSA mechanism. Furthermore, our TSFA retains the data characteristics of the HSI three-dimensional cube through a grouped operation and employs 3D convolutions to extract tokens with spatial and spectral properties. During the computation of self-attention, the TSFA selectively focuses on the most relevant tokens, effectively disregarding the least relevant information. This approach improves interpretive accuracy while reducing redundant computations.
As illustrated in Fig. 4, we first begin by partitioning the input feature $\mathbf{P}\in\mathbb{R}^{C\times H\times W}$ into $g$ groups, each producing a feature map of size $c\times H\times W$. Then we feed them through a 3D point-wise convolution with a kernel size of $1\texttimes 1\texttimes 1$, followed by a 3D depth-wise convolution with a kernel size of $1\texttimes 3\texttimes 3$. This process generates the corresponding query $\mathbf{Q}$, key $\mathbf{K}$, and value $\mathbf{V}$. To facilitate the calculation of self-attention, we reshape and transpose the tensors $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$, adjusting their dimensions by flattening the operation, so each has a shape of $(H\times W\times g)\times c$. These processes can be extended to the multi-head scenario and expressed as follows: